# Advancing African-Accented English Speech Recognition: Epistemic Uncertainty-Driven Data Selection for Generalizable ASR Models

### v3

@@ Abstract
Accents are crucial in human communication as they help us understand others and allow us to communicate intelligibly in a way others understand us. While there has been significant progress in ASR, African-accented ASR has been understudied due to a lack of training datasets which are often expensive to create and demand colossal human labor. Our study aims to address this problem by automating the annotation process and reducing annotation-related expenses through informative uncertainty-based data selection. We propose a new multi-rounds adaptation process that uses epistemic uncertainty and evaluate it across several domains, datasets, and high-performing ASR models. Our results show that our approach leads to a 69.44% WER improvement while requiring on average 45% less data than established baselines. Our approach also improves out-of-distribution generalization for very low-resource accents, demonstrating its viability for building generalizable ASR models in the context of accented African ASR. Moreover, the results of our active learning experiments, simulating real-world settings, where there are no gold transcriptions available, also demonstrate the ability of our approach to favor good quality real-life transcriptions. This indicates that our proposed approach addresses the immediate issue of African-accented ASR and has broader implications for improving ASR systems for other underrepresented and low-resource languages and accents. We open-source the code here.
Advancing African-Accented English Speech Recognition: Epistemic Uncertainty-Driven Data Selection for Generalizable ASR Models
Bonaventure F. P. Dossou1,2,3,4∗, Atnafu Lambebo Tonja5,∗, Chris Chinenye Emezue1,4,6,∗, Tobi Olatunji7,8,∗ Naome A Etori9,∗, Salomey Osei10,∗, Tosin Adewumi11,∗, Sahib Singh12,∗ ∗Masakhane NLP, 1Mila Quebec AI Institute, 2 McGill University, 3 Lelapa AI, 4 Lanfrica 5Instituto Politécnico Nacional, 6Technical University of Munich, 7Intron Health Inc, 8Georgia Institute of Technology 9University of Minnesota - Twin Cities, 10Universidad de Deusto, 11 Luleå University of Technology, 12 Ford Motor Company
@@ 1 Introduction
Generally speaking, Accented English is a form of spoken English that exhibits intonations and pronunciations different from standard English. For instance, Nigerian English differs from British English, which differs from American English. In this work, we consider African-accented English as a form of English spoken by African-native speakers.
Automatic Speech Recognition (ASR) is an active area of research Kodish-Wachs et al. (2018); Finley et al. (2018); Zapata and Kirkedal (2015). In the African context, rich in linguistic diversity and complexity, ASR systems have great potential to reduce language barriers in daily communication. In particular, there is a considerable interest in the development of ASR systems that can accurately distinguish African-accented speech, with a focus on creating systems that can accurately transcribe and interpret speech in a variety of accents and dialects (Koenecke et al., 2020; Hirayama et al., 2012; Tadimeti et al., 2022; Ali et al., 2015; Besacier et al., 2014; Yemmene and Besacier, 2019; Gauthier et al., 2016; Reitmaier et al., 2022; Afonja et al., 2021b). However, developing and deploying ASR systems in an African setting is a challenging task Yemmene and Besacier (2019); De Vries et al. (2014) due to the diversity of African languages, with several dialects and variants characterized by low-resource environments and limited data availability Badenhorst and De Wet (2019, 2017); Barnard et al. (2009); Yemmene and Besacier (2019); DiChristofano et al. (2022). Our methodology facilitates the development of robust, generalizable, and cost-efficient African-accented ASR models. The enhanced generalization for out-of-distribution (OOD) cases, particularly for accents with minimal resources, suggests that our approach extends beyond addressing the specific challenges of African-accented ASR. It holds promise for enhancing ASR systems for other marginalized and resource-scarce languages and accents. This advancement could be a crucial step towards establishing more inclusive and globally accessible ASR technologies. In this work, we reduce this gap, through several contributions listed as follows:
using epistemic uncertainty, we propose a novel multi-rounds adaptation process of several high-performing pretrained ASR models,
we evaluate our approach across several speech domains (clinical, general, both), African-accented speech datasets: AfriSpeech-200 Olatunji et al. (2023), SautiDB Afonja et al. (2021b), MedicalSpeech and CommonVoices English Accented Dataset Ardila et al. (2019), and provide domain-specific as well as accent-specific analyses,
@@
we show that our approach improves the WER performance by 69.44% while reducing significantly the required amount of labeled data (by 35-45%),
@@
we show, based on additional active learning experiments, that our approach is also efficient in real-world settings where there are no gold transcriptions.
2 Challenges and Related Works for African-accented ASR
Although the 2000+ African Languages make up about 31% of the 7000+ languages spoken in the world Eberhard et al. (2019), almost all African languages are classified as "low-resource" Tsvetkov (2017). Despite advancements in ASR over the recent years Kodish-Wachs et al. (2018); Blackley et al. (2019), research or applications in the African context are not well investigated in the literature, likely a product of the paucity of data for accented languages, inadequate research funding, data-hungry approaches well-suited to high-resource settings, and most importantly, data-efficient approaches are inadequately explored. As a result of scarce data and resources, it is also challenging to pretrain large multilingual language models for African languages Dossou et al. (2022).
Despite its potential to improve language communication, ASR models suffer from racial bias and poor performance on accented speech Koenecke et al. (2020), particularly African accents. Recognizing heavy accents in speech data is a challenge for ASR systems since it contains high variability in terms of pronunciations Das et al. (2021). In recent years, we have seen a drastic improvement in ASR systems due to more advanced deep learning systems and the availability of data. It is however shocking that racial disparities and poor performances in accented speech still exist for black speakers. Typically, ASR system generalization is hard to achieve due to difficulties in getting good speech transcript data, including for African accents. The different dialects and accents of many languages introduce some level of difficulty in transcribing such data. These variations are much more complicated in the case of African languages, as speech data in general is more difficult to source. For this reason, there is a poor performance of ASR in African languages.
Improving ASR on African accented speech and linguistic diversity is hence necessary to avoid speakers having to change their way of speaking Mengesha et al. (2021) (a situation that will likely make speakers feel excluded, and the social impact is high), while bringing enormous potential to transform several working areas and daily human needs and communication.
@@ 3 Datasets and Methodology
In this work, we established several baselines, across several high-performing ASR systems. Our baselines are produced by feeding the pretrained ASR models, the entirety (100%) of available labeled data. To investigate the effectiveness of our approach, we impose data constraints, choosing fewer training examples than available, to not exceed 60-65% of the entire initial dataset. This simulates a more realistic scenario where the entirety of the data might not be available or feasible to use, thus testing the robustness and efficiency of our method under constrained conditions. We utilize 30% of the initial dataset as our training set and set the remaining 70% as the pool dataset. The purpose of the pool dataset in our study is to allow the model, in our proposed framework, to learn to carefully select the most informative data samples.
To further solidify and demonstrate the effectiveness of our method, we leveraged it in an active learning (AL) setup. AL is a semi-supervised machine learning algorithm that makes use of only a few initial training data points to achieve better performance of a given model. It is important to note that due to the lack of African-accented ASR benchmarks, we primarily compared the results, to our baselines and our best-performing selection technique. Within the defined constraints, for our experiments, for our proposed method, we adjust the sizes of the training and pool datasets, based on the number of training examples available for each domain (clinical, general, and both) (see Tables 2, 4 and Appendix A).
@@ 3.1 Datasets
In this work, we primarily used the AfriSpeech-200 dataset Olatunji et al. (2023), a novel 200-hour African-accented English speech corpus carefully curated for clinical and general domains ASR. The dataset covers 120+ African accents, all derived from 5 language families Eberhard et al. (2019): Afro-Asiatic, Indo-European, Khoe-Kwadi (Hainum), Niger-Congo, and Nilo-Saharan, representing regional diversity across western, eastern, and southern Africa. AfriSpeech-200 is a crowd-sourced dataset involving over 2000 African speakers, representing 13 anglophone countries across sub-Saharan Africa and the US (see Table 1).
To establish the dataset-agnostic nature of our approach, we explore three additional datasets:
1.
SautiDB Afonja et al. (2021a): a dataset of Nigerian accent recordings with 919 audio samples, a sampling rate of 48kHz each, for a total amount of 59 minutes of recordings,
2.
MedicalSpeech111https://www.kaggle.com/datasets/paultimothymooney/medical-speech-transcription-and-intent: a dataset which contains 6,661 audio utterances for common medical symptoms like knee pain or headache, for a total of 8 hours of recordings;
3.
CommonVoices English Accented Dataset: a subset of English CommonVoices (version 10) Ardila et al. (2019), with western accents removed as our primary interest is in low-resource settings.
@@
| Item | Train (*AL Train) | Dev | Test | AL Top-k |
@@ 3.2 Methodology
Our methodology is based on epistemic uncertainty. Epistemic Uncertainty (EU) is the uncertainty of the model, most often due to a lack of training data. For a given input , EU is usually defined with the variance (for a given model with some parameters ) as follows:
where . is the learned approximation of . In our experiments, we decided to use the standard deviation defined (STD) as . EU is useful for the model’s training as it improves its robustness and encourages exploration and diversity, to mitigate inductive bias related to underrepresented accents (accents with fewer speech data and speakers).
One of the most popular ways to estimate uncertainty is by inferring predictive distributions with Bayesian Neural Networks (BNNs). A predictive distribution is defined as where is the target, is the input and is the dataset. One can thus inspect the variance (or the standard deviation, as we used in our experiments) and uncover uncertainty, using the predictive distribution. A way of learning the predictive distribution is to learn a distribution over the parameters i.e. the parametric posterior distribution .
There are many ways of computing epistemic uncertainty, but the two most common approaches are Monte Carlo Dropout Gal and Ghahramani (2016) and Deep Ensembles Lakshminarayanan et al. (2017). Many empirical results demonstrated that Deep Ensembles generally work better while being very computationally expensive. Due to computational resource constraints, we decided to demonstrate the effectiveness of our approach using Mc-Dropout empirically. Mc-Dropout provides a scalable way to learn a predictive distribution. It randomly switches off neurons acting as regularizers in a given neural network. Each dropout configuration corresponds to a different sample from the approximate parametric posterior distribution . Sampling from the (approximate) posterior enables Monte Carlo integration of the likelihood of the model, which uncovers the predictive distribution. Our uncertainty-based selection pipeline is detailed in Figure 1 and the algorithm 1. In each adaptation round, we use a finetuned model, to select the most uncertain samples from the pool dataset and add them to the training set for the next round of the model adaptation.
While performing AL experiments, we consider (mask) the samples from the pool dataset to be entirely unlabeled (i.e. no gold transcriptions as targets). This implies that we simulate an unsupervised context, and learn to select the best possible predicted transcription as the target (gold) transcriptions.
Using Mc-Dropout sets a pre-requisite for the model to be equipped with Dropout components during pretraining. This condition eliminates some high-level ASR models like Whisper whisper. However, we finetuned Whisper in our baseline setting and evaluated it against our best-performing approach. This is to assess its robustness and adequacy. In this work, we use the following four state-of-the-art (SOTA) ASR models: Wav2Vec2-XLSR-53 Conneau et al. (2020), Hubert-Large Hsu et al. (2021), WavLM-Large Chen et al. (2022), and NVIDIA Conformer-CTC Large (en-US) Gulati et al. (2020), which we hereafter refer to as Wav2Vec, Hubert, WavLM, and Nemo respectively.
@@ 3.3 Experimental Design
Added to the baselines described above, in our framework, we explored two selection techniques: most and random. The most technique consists of selecting the most uncertain audio samples from the pool dataset, while the random one, consists of selecting random audio samples across the pool dataset. In this work, we used the words "sampling", "techniques", and "selection" interchangeably with the same meaning.
We performed various experiments to evaluate our proposed approach’s efficiency and applicability. Each experiment at the model level is performed in three domains: general, clinical, and both (which is the combination of general and clinical data).
For each domain, we have four setups: (a) baseline which consists of simply finetuning a pretrained model on the entire dataset (as described in section 3.1), (b) EU-Most where the pretrained model is finetuned within our framework using the most selection, (c) Random uses the random selection, and (d) AL (w/ EU-Most) where the pretrained model is fine-tuned using active learning and the most technique.
Our experiments are designed to allow us to answer the following specific questions:
@@
given a set of African accents , how does the pretrained ASR model adapt to across the adaptation rounds and domains (clinical, general and both)?
@@
which selection strategy between most and random works better, and for which domain(s)?
@@
which domain(s) help(s) the model to do better, and how does the model perform (in terms of uncertainty) across the domain(s)?
@@
is epistemic uncertainty-based selection effective in low-resource data scenarios?
@@
To answer questions 1-4, we define a metric called Uncertainty WER (U-WER). U-WER is described as follows:
where is the speech input, the gold target (transcript), the finetuned ASR model. Due to the nature of the MC-Dropout technique, running it times will produce different speech representations for the original speech input . Each speech representation is denoted , which could consequently lead to different predicted transcription, that we denote by . is the WER function, computing the WER between the predicted transcript and the gold (true) transcript . is the average of WERs between the different speech representations and the target prediction. In our experiments, . The ideal case would be when U-WER0, because this would mean that the model is robust enough to produce consistent transcripts and speech representation instances .
The rationale behind the U-WER is the following: as we acquire new data points that are intrinsically beneficial to the model and its learning, the U-WER should (consistently) decrease or be constant across the multiple adaptation rounds. We could therefore affirm that the model is gaining in robustness, knowledge, and performance; which is a very important aspect of generalization. This last point will be enforced by answering question 5 for which we used the standard WER.
To answer question 5, we evaluated our approach with three additional ASR pretrained models (Nemo, WavLM, and Hubert), and across 3 external datasets (SautiDB, Common Voice English Accented, and MedicalSpeech). Additionally, to answer questions 1-4, for consistency and better visualization, we consider exclusively the top-10 (in terms of frequency) accents present across three adaptation rounds and both selection methods. Regarding the very low-resource settings, we consider the 5 accents with the least number of hours of recordings.
In our AL setup, the U-WER is computed using pairwise WER scores among the predicted transcriptions, rather than using the gold transcriptions. To select the ideal or best predicted transcript for the unlabeled speech , we proceed as described in Algorithm 2.
This implies that for the current speech input , is computed by taking the standard deviation ‘wer_list‘ variable (this corresponds exactly to the formula of the described above). Finally, we assign to , the predicted transcription with the lowest average WER i.e. the lowest value of the dictionary ‘wer_target_dict‘. For our experiments, we use various types of GPUs: 6 RTX 800 (with 48 and 128 GB), and 4 A100 (with 48 and 128 GB). Our training and evaluation were constantly deployed over a month. Our models have 311281844 trainable parameters. Each audio sample has been normalized and processed with a sample rate of 16kHz.
| Model | General | Clinical | Both | |||||||||
| Baseline | Random | EU-Most | AL (w/ EU-Most) | Baseline | Random | EU-Most | AL (w/ EU-Most) | Baseline | Random | EU-Most | AL (w/ EU-Most) | |
| Wav2vec | 0.4980 | 0.1111 | 0.1011 | 0.1059 | 0.5610 | 0.3571 | 0.2457 | 0.2545 | 0.5300 | 0.1666 | 0.1266 | 0.1309 |
| **Hubert | 0.1743 | ✗ | 0.1901 | 0.1887 | 0.2907 | ✗ | 0.2594 | 0.2709 | 0.2365 | ✗ | 0.2453 | 0.2586 |
| **WavLM | 0.1635 | ✗ | 0.1576 | 0.1764 | 0.3076 | ✗ | 0.2313 | 0.2537 | 0.2047 | ✗ | 0.1897 | 0.1976 |
| *Nemo | 0.2824 | ✗ | 0.1765 | 0.1815 | 0.2600 | ✗ | 0.2492 | 0.2526 | 0.3765 | ✗ | 0.2576 | 0.2610 |
| Average Performance | 0.2796 | - | 0.1563 | 0.1631 | 0.3548 | - | 0.2464 | 0.2579 | 0.3369 | - | 0.2043 | 0.2120 |
| Whisper-Medium Olatunji et al. (2023) | 0.2806 | - | - | - | 0.3443 | - | - | - | 0.3116 | - | - | - |
| Dataset | Split and Size for our approach | Finetuning Epochs | Baseline | Uncertainty Sampling w/ most | |||
| Train | Aug | Top-k | Test | (Entire training dataset) | (Train + Aug) | ||
| SautiDB Afonja et al. (2021a) | 234 | 547 | 92 | 138 | 50 | 0.50 | 0.12 |
| MedicalSpeech | 1598 | 3730 | 1333 | 622 | 5 | 0.30 | 0.28 |
| CommonVoices English Accented Dataset (v10.0) Ardila et al. (2019) | 26614 | 62100 | 10350 | 232 | 5 | 0.50 | 0.22 |
| Average | ✗ | ✗ | ✗ | ✗ | ✗ | 0.43 | 0.20 |

### v4

@@ Abstract
Accents play a pivotal role in shaping human communication, enhancing our ability to convey and comprehend messages with clarity and cultural nuance. While there has been significant progress in Automatic Speech Recognition (ASR), African-accented English ASR has been understudied due to a lack of training datasets, which are often expensive to create and demand colossal human labor. Combining several active learning paradigms and the core-set approach, we propose a new multi-rounds adaptation process that uses epistemic uncertainty to automate the annotation process, significantly reducing the associated costs and human labor. This novel method streamlines data annotation and strategically selects data samples that contribute most to model uncertainty, thereby enhancing training efficiency. We define a new metric called U-WER to track model adaptation to hard accents. We evaluate our approach across several domains, datasets, and high-performing speech models. Our results show that our approach leads to a 69.44% WER improvement while requiring on average 45% less data than established baselines. Our approach also improves out-of-distribution generalization for very low-resource accents, demonstrating its viability for building generalizable ASR models in the context of accented African ASR. We open-source the code here.
@@ 1 Introduction
Automatic Speech Recognition (ASR) is an active research area that powers voice assistant systems (VASs) like Siri and Cortana, enhancing daily communication [27, 17, 46]. Despite this progress, no current VASs include African languages, which account for about 31% of the world languages, and their unique accents [16, 44]. This gap underscores the need for ASR systems that can handle the linguistic diversity and complexity of African languages, especially in crucial applications like healthcare. Due to the lack of representation of these languages and accents in training data, existing ASR systems often perform inadequately, even mispronouncing African names [36].
To address these challenges, our work focuses on adapting pretrained speech models to better transcribe African-accented English, defined by unique intonations and pronunciations [7, 22]. We use epistemic uncertainty (EU) [25] to guide the adaptation process by identifying gaps in model knowledge and prioritizing data for the model to learn from next. This is particularly beneficial in scenarios where data annotation is costly or time-consuming, as often seen in the African context [5, 4, 6, 45, 12, 14, 13]. EU also improves robustness and encourages exploration to mitigate inductive bias from underrepresented accents. Common approaches to compute EU include Monte Carlo Dropout (Mc-Dropout) [18] and Deep Ensembles [30], with the latter being more effective but computationally expensive. Due to resource constraints, we use Mc-Dropout, which requires models to have dropout components during pretraining.
Active Learning (AL) is a machine learning (ML) technique that leverages epistemic uncertainty (EU) to select the most informative data points from an unlabeled dataset for labeling, thereby improving model performance with fewer training instances. Common types of AL include Deep Bayesian Active Learning (DBAL) [19, 23] and Adversarial Active Learning (AAL) [15]. AAL selects examples likely to be misclassified by the current model, refining it iteratively by challenging it with difficult cases to enhance robustness. The core-set approach (CSA) [42] is also related, as it selects a subset of training data to ensure a model trained on this subset performs comparably to one trained on the full dataset, addressing scalability and efficiency.
Acquisition functions (AFs) in AL select the most informative samples from an unlabeled dataset for labeling. Key AFs include uncertainty sampling (US) [31], Bayesian Active Learning by Disagreement (BALD) [19], and BatchBALD [26]. US targets data points with the highest model uncertainty. BALD maximizes the mutual information between model parameters and predictions. BatchBALD is an extension of BALD which selects multiple samples simultaneously but may choose redundant points. US is the least computationally expensive, ideal for efficient data labeling.
In this work, we leverage and combine DBAL, AAL, US, and CSA in the following way (in order):
1.
we integrate the CSA by leveraging smaller training subsets ( smaller than the full available training sets)
- 2.
3.
we use the estimated EU, and integrate the idea of AAL by using the US acquisition function.
We evaluate our approach across several domains (general, clinical, general+clinical aka both), several datasets (AfriSpeech-200 [37], SautiDB [2], MedicalSpeech, CommonVoices English Accented Dataset [3]), and several high-performing speech models (Wav2Vec2-XLSR-53 [10], HuBERT-Large [24], WavLM-Large [8], and NVIDIA Conformer-CTC Large (en-US) [20]). Our results show a 69.44% WER improvement while requiring on average 45% less data than established baselines. We also adapt the standard Word Error Rate (WER) to create a new metric called Uncertainty WER (U-WER) to track model adaptation to African accents.
The impact of our approach is substantial. It develops more robust, generalizable, and cost-efficient African-accented English ASR models and reduces dependency on large labeled datasets, enabling deployment in various real-world scenarios. Our results show improved generalization for out-of-distribution (OOD) cases, especially for accents with minimal resources, addressing specific challenges in African-accented ASR. Additionally, by focusing on equitable representation in ASR training, our methodology promotes fairness in AI, ensuring technology serves users across diverse linguistic backgrounds without bias [41, 34, 32]. Our contributions are listed as follows:
we combine DBAL, AAL, CSA, and EU to propose a novel way to adapt several high-performing pretrained speech models to build efficient African-accented English ASR models,
we define a new and simple metric called U-WER that allows us to measure and track how the variance of the model, across hard accents, changes over the adaptation process,
@@
we show that our approach improves the WER performance by 69.44% while significantly reducing the required amount of labeled data (by 45
@@
we show, based on additional AL experiments, that our approach is also efficient in real-world settings where there are no gold transcriptions.
2 Background and Related Works
2.1 Challenges for African-accented ASR
State-of-the-art (SOTA) ASR technologies, powered by deep learning and neural network architectures like transformers, achieve high accuracy with Standard American English and major European languages. However, they often fail with African accents due to high variability in pronunciation and lack of quality speech data [29, 11]. This results in racial bias, poor performance, and potential social exclusion as speakers might alter their speech to be understood [29, 28, 9, 33]. Enhancing ASR for African languages is crucial for equitable voice recognition, especially in healthcare, education, and customer service. Solutions should focus on diversifying training datasets and developing robust modeling techniques tailored to these languages unique characteristics.
2.2 Active Learning
AL aims at reducing the number of training examples to be labeled by automatically processing the unlabeled examples and then selecting the most informative ones concerning a given cost function for a human to label. It is particularly effective when labeled data is scarce or expensive, optimizing the learning process by focusing on samples that most improve the model performance and generalization [43, 19]. Several works have demonstrated its effectiveness and efficiency. An AL setup involves an unlabeled dataset , a labeled training set , and a predictive model with likelihood parameterized by . The setup assumes the presence of an oracle to provide predictions for all . After training, a batch of data is selected from based on its EU.
In [21], AL was applied to a toy dataset of How May I Help You recordings. Confidence scores were estimated for each word and used to compute the overall confidence score for the audio sample. This approach achieved competitive results using 27% less data compared to the baseline. In [40], the authors estimated confidence scores for each utterance using an online algorithm with the lattice output of a speech recognizer. The utterance scores were filtered through an informativeness function to select an optimal subset of training samples, reducing the labeled data needed for a given WER by over 60%. [35] experimented with AL for accent adaptation in speech recognition. They adapted a source recognizer to the target accent by selecting a small, matched subset of utterances from a large, untranscribed, multi-accented corpus for human transcription. They used a cross-entropy-based relevance measure alongside uncertainty-based sampling. However, their experiments on Arabic and English accents showed worse performance compared to baselines, while using more hours of recordings.
@@ 3.1 Datasets
We used the AfriSpeech-200 dataset [37], a 200-hour African-accented English speech corpus for clinical and general ASR. This dataset includes 120+ African accents from five language families: Afro-Asiatic, Indo-European, Khoe-Kwadi (Hainum), Niger-Congo, and Nilo-Saharan, representing regional diversity across Africa. It was crowd-sourced from over 2000 African speakers from 13 anglophone countries in sub-Saharan Africa and the US.
To demonstrate the dataset-agnostic nature of our approach, we also explored three additional datasets: (1) SautiDB [1], Nigerian accent recordings with 919 audio samples at a 48kHz sampling rate, totaling 59 minutes; (2) MedicalSpeech111https://www.kaggle.com/datasets/paultimothymooney/medical-speech-transcription-and-intent, containing 6,661 audio utterances of common medical symptoms, totaling 8 hours; and (3) CommonVoices English Accented Dataset, a subset of English Common Voice (version 10) [3], excluding western accents to focus on low-resource settings.
@@
| Item | Train ( Train) | Dev | Test | AL Top-k |
@@ 3.2 Methodology
In our approach, to compute EU for a given input , we perform Mc-Dropout to obtain multiple stochastic forward passes through a finetuned ASR model with likelihood where is the weights of . Let be a function that computes the WER between the predicted and the target transcripts. Let be the number of stochastic forward passes. For each pass , we apply dropout, obtain the output transcript, and compute the WER:
where is a binary mask matrix sampled independently for each pass. EU(()) can then be estimated from the stochastic forward passes as follows:
The use of Mc-Dropout requires models to have dropout components during training. This excludes some models like Whisper [38], which we still finetuned and evaluated as a baseline. We use four state-of-the-art pretrained models: Wav2Vec2-XLSR-53, HuBERT-Large, WavLM-Large, and NVIDIA Conformer-CTC Large (en-US), referred to as Wav2Vec, Hubert, WavLM, and Nemo, respectively.
3.2.1 Uncertainty WER
To handle diverse accents, we aim to reduce the EU of the models across hard accents after each adaptation round. We define a metric called U-WER to track this. To compute U-WER() where is a hard accent, we condition EU on :
where is the audio sample with accent a and
Ideally, U-WER0. The rationale behind U-WER is that as beneficial data points are acquired, U-WER should decrease or remain constant, indicating increased robustness, knowledge, and performance, crucial for generalization. During AL, U-WER is computed using pairwise WER scores among predicted transcriptions, not gold transcriptions (see section 3.3). To select the best-generated transcript for unlabeled speech , we follow Algorithm 2.
@@ 3.3 Experimental Design
To work within our framework, we define the following selection strategies:
random: Randomly selects audio samples from .
EU-Most: Selects the most uncertain audio samples from to add to .
AL-EU-Most: Combines AL with the EU-Most strategy to finetune the pretrained model.
We also define standard fine-tuning (SFT) as baseline using all available data for finetuning. In SFT, is empty. While running the defined strategies in our framework, we impose data constraints, not exceeding 60-65% of the initial dataset after all adaptation rounds. is 30% of , and is 70% of . This simulates realistic scenarios where not all data might be available, testing the approach’s robustness and efficiency under constraints. The number of samples in and is based on available training examples for each domain (see Tables 2, 4, and Appendix A.1).
Our EU-based pipeline is shown in Figure 1 and Algorithm 1. In each adaptation round, we use a finetuned model and a selection strategy to choose samples from to add to . During AL experiments, we consider samples from as unlabeled: (1) using Mc-Dropout, we obtain different input representations per audio sample to get different transcripts; (2) we then learn to select the best-generated transcription as the target transcription according to Algorithm 2.
Our experiments aim to answer the following research questions:
@@
how does the pretrained ASR model adapt to a set of African accents across adaptation rounds and domains?
@@
which selection strategy (EU-most or random) works better, and for which domain(s)?
@@
which domain(s) help the model perform better, and how does the model perform (in terms of uncertainty) across the domain(s)?
@@
what is the impact of EU-based selection on the model’s efficiency in low-resource data scenarios?
@@
U-WER will answer question 4. To answer question 5, we evaluated our approach with three additional pretrained models (Nemo, WavLM, and Hubert) and across three external datasets (SautiDB, CommonVoices English Accented Dataset, and MedicalSpeech). For consistency and better visualization, we considered the top-10 (in terms of frequency) accents across three adaptation rounds and both selection strategies to answer questions 1-4. For very low-resource settings, we considered the five accents with the least recording hours.
We used various GPUs for our experiments: 6 RTX 800, and 4 A100. Training and evaluation were conducted over a month. Our models have approximately 311 million trainable parameters. Each audio sample was normalized and processed at a 16kHz sample rate. We used default parameters from the HuggingFace library for each pretrained model.
| Model | General | Clinical | Both | ||||||
| Baseline | EU-Most | AL-EU-Most | Baseline | EU-Most | AL-EU-Most | Baseline | EU-Most | AL-EU-Most | |
| Wav2vec | 0.4980 | 0.1011 | 0.1059 | 0.5610 | 0.2457 | 0.2545 | 0.5300 | 0.1266 | 0.1309 |
| **Hubert | 0.1743 | 0.1901 | 0.1887 | 0.2907 | 0.2594 | 0.2709 | 0.2365 | 0.2453 | 0.2586 |
| **WavLM | 0.1635 | 0.1576 | 0.1764 | 0.3076 | 0.2313 | 0.2537 | 0.2047 | 0.1897 | 0.1976 |
| *Nemo | 0.2824 | 0.1765 | 0.1815 | 0.2600 | 0.2492 | 0.2526 | 0.3765 | 0.2576 | 0.2610 |
| Average Performance | 0.2796 | 0.1563 | 0.1631 | 0.3548 | 0.2464 | 0.2579 | 0.3369 | 0.2043 | 0.2120 |
| Whisper-Medium | 0.2806 | - | - | 0.3443 | - | - | 0.3116 | - | - |